OpenStudy (anonymous):

find the value of x logx+log(2x-1)=0

OpenStudy (anonymous):

log(x) + log(2x-1) = 0 log (x*(2x-1))= 0 10^0 = 2x^2 - x 2x^2 - x -1 =0 (2x+1)(x-1) = 0 x = -1/2 or 1 x cannot be -1/2 (because it would cause log(2x-1) = log(0) Therefore, x = 1

OpenStudy (anonymous):

thank you